# Robust Adaptive Time-Varying Control Barrier Function with Application to Robotic Surface Treatment

### v1

@@ I Introduction
Set invariance can be achieved in many ways, and using Control Barrier Functions (CBFs) [6] is one of the promising approaches, ensuring forward invariance of a given set. CBFs can be easily implemented with quadratic programming and have been widely used in robotics applications [7][9][10]. Although CBFs accomplish the forward invariance of a given static constraint set, they are not applicable to applications where a constraint set is changed over time: a time-varying set. In this vein, using Time-Varying Control Barrier Functions (TVCBFs) is one of the potential ways to enforce time-varying constraints and successfully applied to human-assist wheelchair control [15] and dynamic obstacle avoidance [16].
However, since CBF-based controllers are highly dependent on the model accuracy, the forward invariance of a given set might not be satisfied if the model uncertainties such as parametric uncertainty and input disturbances are not taken into account. To resolve this issue, adaptive and robust control schemes have been recently proposed. Robust adaptive CBFs (RaCBFs) [12] introduce the tightened safe set by imposing robustness against the uncertainty and can be integrated with Set Membership IDentification (SMID) [13] to estimate the maximum bound of the uncertainty, so that less conservatism can be achieved. Another uncertainty, input disturbances are frequent in real-world implementations; these deteriorate the performance of CBFs-based controllers. The notion of Input-to-State Safety (ISSf) is defined to treat input disturbances, making a given safe set inflated based on the magnitude of the disturbances [14]; thus they enable minimizing safety violations. Notably, in [17], TVCBFs are combined with ISSf conditions to impose robustness against the bounded input disturbances. However, they do not consider inherent model uncertainty in the system model like RaCBFs, so the violation of constraints could occur.
Figure 1: Block diagram of the control system including a robust adaptive time-varying control barrier function with a given time-varying safe set, $\mathcal{S}(t)$. Given an uncertain system with parametric uncertainty and additive disturbances, the safety filter ensures that the system states stay in a robust subset, $\mathcal{S}_{\bm{\theta}_{\delta}}^{r}\subseteq\mathcal{S}(t)$. Data-driven update method, Set Membership IDentification (SMID) is used to reduce the maximum possible error $||\bm{\tilde{\vartheta}}||$.
@@
We demonstrate the proposed method on robotic surface treatment application in this paper. Generally, the robotic surface treatment can be accomplished by maintaining a constant contact force between the polishing tool and the surface of a given workpiece [1]. Many force control schemes including passive [2][3] and active compliance controllers [4] have been used in robotic surface treatment.
Even if the current robotic systems for surface treatment have focused on controlling constant contact forces to achieve an acceptable quality, they often neglect contact pressures over the surface, which leads to failure to keep uniform quality. Moreover, formal quality guarantees have not been fully investigated yet, which is a main challenge in robotic surface treatment applications [YkCASE2022][23]. To remedy this challenge, the proposed method is suitable for the robotic surface treatment application to accomplish quality guarantees based on a set invariance scheme. This is achieved by the fact that the proposed method regulates the time-varying reference contact force trajectory, which in turn governs the contact pressure, and then we ultimately regulate the quality of robotic surface treatment.
@@ II Motivating Example
In this section, we describe an application example that motivates this work. The proposed method is inspired by a robotic surface treatment task where the quality can be determined based on the material removal rate (MRR) [1]. The MRR is modeled using Preston’s equation [18]:
@@ III-A Time-Varying Control Barrier Functions
where $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state, $f,g$ are locally Lipschitz continuous functions, $\bm{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ represents a control input. Consider a time-varying safety set $\mathcal{S}$ defined as $\mathcal{S}=\{\bm{x}\in\mathcal{X}\text{ }|\text{ }h(\bm{x},t)\geq 0\}$. A continuously differentiable function $h:\mathcal{X}\times\mathbb{R}\rightarrow\mathbb{R}$ is a Time-Varying Control Barrier Function (TVCBF) if there exist constants $K>0$ and $C>0$ satisfying the following for all $t\geq 0$ [15]:
@@ III-B Input-to-State Safety Control Barrier Functions
Unmodeled dynamics and disturbances may result in safety violations if we directly use standard CBFs [6]. We now introduce the input disturbance, $\bm{d}\in\mathbb{R}^{m},||\bm{d}||_{\infty}\leq\delta,\delta\geq 0$ in (2):
@@
where $\mathcal{S}\subseteq\mathcal{S}_{\delta}$ and $h_{\delta}(\bm{x})=h(\bm{x})+\gamma(\delta)$. We say that the system (5) is Input-to-State Safe (ISSf) if there exists $\gamma$ and $\delta$ such that for all $\bm{d}(t)$, $\mathcal{S}_{\delta}$ is forward invariant [14]. An ISSf control barrier function is defined as follows:
Theorem 1 ([14])
@@ III-C Robust adaptive Control Barrier Functions (RaCBFs)
In the presence of parametric uncertainty in the system, we consider the following control affine system and adaptive parameter estimator dynamics simultaneously [11]:
$\begin{bmatrix}\dot{\bm{x}}\\ \dot{\hat{\bm{\theta}}}\end{bmatrix}=\begin{bmatrix}f(\bm{x})+F(\bm{x})\bm{\theta}^{*}+g(\bm{x})\bm{u}\\ \Gamma\tau(\bm{x},\hat{\bm{\theta}})\end{bmatrix}$ (8)
@@
$\lambda_{\text{min}}(\Gamma)\geq\frac{||\tilde{\bm{\vartheta}}||^{2}}{2h_{r}(\bm{x},\bm{\theta})}.$
@@ III-D Set Membership IDentification (SMID)
Set Membership IDentification (SMID) is leveraged to compute a bound on the parameter estimation error based on input-output data [13]. We consider a linear regression equation
@@
where $d_{j}$ is a disturbance, $\bm{\theta}$ is a vector of unknown constant parameters, and $\mathcal{Y}_{j}$, $\mathcal{D}_{j}$ are measured quantities. Let =$\mathcal{H}=\{\mathcal{Y}_{j},\mathcal{D}_{j}\}^{M}_{j=1}$ be a data history with size $M\in\mathbb{N}$. To bound the value of $\bm{\theta}^{*}$ in (8), we define $\mathcal{Y}_{j}=\dot{\bm{x}}(t_{j})-f(\bm{x}(t_{j}))-g(\bm{x}(t_{j}))\bm{u}(t_{j})$ and $\mathcal{D}_{j}=F(x(t_{j}))$, where $t_{j}$ is the time at which the data is recorded. Consider the following sequence of bounded parameter sets with $\Xi_{0}=[\underline{\bm{\theta}}_{0},\overline{\bm{\theta}}_{0}]$ and $\forall j\in M$:
@@
where $[\underline{\bm{\theta}}_{0},\overline{\bm{\theta}}_{0}]$ is the initial lower and upper bound of the unknown parameter, and $\bm{1}_{n}$ is an $n-$dimensional vector of ones, and $\varepsilon\in\mathbb{R}$ is the precision variable that must be chosen such that $||d_{j}||_{\infty}\leq\epsilon$ for all $j$. The sequence of sets, $\Xi_{t_{j}}$ can be calculated based on Linear Programming (LP) such that $\Xi_{t_{j}}\subseteq\Xi_{t_{j-1}}\subseteq\cdots\subseteq\Xi_{0}$ and $\bm{\theta}\in\Xi_{t_{j}}$ [20].
@@ IV-A Parametric Adaptive System Model with Uncertainties
where $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state, $f,g$ are locally Lipschitz continuous functions, $\bm{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ represents a control input, and $\bm{d}(t)\in\mathbb{R}^{m}$ is a bounded input disturbance, $||\bm{d}(t)||_{\infty}\leq\delta\in\mathbb{R}_{+}$, and $F:\mathcal{X}\rightarrow\mathbb{R}^{n\times k}$ is smooth on $\mathcal{X}$ and models parametric uncertainty for $f$ such that $F(\bm{0})=\bm{0}$, and $\bm{\theta}^{*}\in\Theta\subset\mathbb{R}^{k}$ is the vector of unknown parameters. In the following, we include the time derivative term of $h_{r}$ to deal with time-varying constraints.
@@ IV-B Robust Adaptive Time-Varying Control Barrier Function
where $\gamma(\cdot)$ is the extended class $\mathcal{K}_{\infty}$ function. And if we choose $\lambda_{\text{min}}(\Gamma)\geq\frac{||\tilde{\bm{\vartheta}}||^{2}}{2h_{r}(\bm{x},\bm{\theta},t)}$ to satisfy $h=h_{r}-\frac{1}{2}\tilde{\bm{\theta}}^{\top}\Gamma^{-1}\tilde{\bm{\theta}}\geq h_{r}-\frac{1}{2}\tilde{\bm{\vartheta}}^{\top}\Gamma^{-1}\tilde{\bm{\vartheta}}\geq 0$ according to [12], then we obtain the following definition:
@@
A continuously differentiable function $h$ is a Robust adaptive Time-Varying Control Barrier Function (RaTVCBF) for the system (15) with bounded disturbance, $||\bm{d}(t)||_{\infty}\leq\delta$ if there exist extended class $\mathcal{K}_{\infty}$ functions $\alpha$ and $\gamma$ such that for any $\bm{\theta}\in\Theta$ and for all $\bm{x}\in\mathcal{S}^{r}_{\bm{\theta}_{\delta}}$:
@@
with $\text{with }\qquad\eta=\Big{|}\Big{|}L_{g}h_{r}(\bm{x},\bm{\theta},t)\Big{|}\Big{|}\delta,$
@@
where $\eta=\Big{|}\Big{|}L_{g}h_{r}(\bm{x},\bm{\theta},t)\Big{|}\Big{|}\delta$, $L_{f}h_{r}=\frac{\partial h_{r}}{\partial\bm{x}}(\bm{x},\hat{\bm{\theta}},t)f(\bm{x})$, $L_{F}h_{r}=\frac{\partial h_{r}}{\partial\bm{x}}(\bm{x},\hat{\bm{\theta}},t)F(\bm{x})$, and $L_{g}h_{r}=\frac{\partial h_{r}}{\partial\bm{x}}(\bm{x},\hat{\bm{\theta}},t)g(\bm{x})$, renders the system (15) ISSf with respect to $\mathcal{S}_{\bm{\theta}_{\delta}}^{r}$.
@@
The proof follows [12]. Let a candidate control barrier function $h$ be:
@@
$\dot{h}(\bm{x},\hat{\bm{\theta}},t)=\dot{h}_{r}(\bm{x},\hat{\bm{\theta}},t)+\tilde{\bm{\theta}}^{\top}\Gamma^{-1}\dot{\hat{\bm{\theta}}}\$ (21)
@@
$\frac{\partial h_{r}}{\partial\bm{x}}F(\bm{x})\Bigg{(}\hat{\bm{\theta}}-\Gamma\bigg{(}\frac{\partial h_{r}}{\partial\hat{\bm{\theta}}}\bigg{)}^{\top}\Bigg{)}.$ (22)
Then, we obtain the following equation and inequalities with $\lambda(\bm{x},\hat{\bm{\theta}},t)\triangleq{\hat{\bm{\theta}}}-\Gamma\big{(}\frac{\partial h_{r}}{\partial\hat{\bm{\theta}}}(\bm{x},\hat{\bm{\theta}},t)\big{)}^{\top}$ from [11]:
@@
$\quad-\Big{|}\Big{|}L_{g}h_{r}\Big{|}\Big{|}\delta+\frac{\partial h_{r}}{\partial t}(\bm{x},\hat{\bm{\theta}},t).$ (23)
In the last inequality of (23), we choose $\tau(\bm{x},\hat{\bm{\theta}},t)=-\Big{(}\frac{\partial h_{r}}{\partial\bm{x}}F(\bm{x})\Big{)}^{\top}$ similar to (9) provided by [11], and then the second term will be zero. Afterwards, we introduce the maximum possible error, $\tilde{\bm{\vartheta}}$, and if the following inequality is satisfied:
@@
where $\eta=\Big{|}\Big{|}L_{g}h_{r}\Big{|}\Big{|}\delta$, and we choose $\gamma(\delta)=-\alpha^{-1}(-\frac{\epsilon\delta^{2}}{4})$ from [21], and then finally the following sufficient condition of forward invariance of the set, $\mathcal{S}^{r}_{\bm{\theta}_{\delta}}$ can be obtained:
@@
[…]
@@ V Application to Robotic Surface Treatment
In this section, it is shown that the proposed method ensures that MRR (1) remains within an acceptable range across the entire surface. According to [1], controlling MRR within the desired range ultimately achieves the desired quality guarantees. We verify the proposed method with numerical simulation on MATLAB and conduct experimental validation in a real robotic setup.
@@ V-A Robotic Surface Treatment
To ensure quality guarantees of the surface treatment via MRR as we mentioned in Section II, we enforce contact force constraints in (25). First, we define the parametric contact force, $\bm{f}_{c}\in\mathbb{R}^{n}$ from the Kelvin-Voigt model [22] as:
@@
Figure 3: Pipeline of the robotic surface treatment. First, a user defines the desired MRR. Subsequently, in the force reference generator, we estimate contact area between a tool and a given surface and then calculate the desired contact pressure based on the desired MRR. Afterwards, the desired force is obtained from the desired contact pressure by using (1). Finally, we can determine force bounds (e.g. ±$\pm 15\%$ of the desired force) that are used as constraints to ensure the desired MRR in the dotted box.

### v2

@@ Abstract
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Set invariance can be achieved in many ways, and using Control Barrier Functions (CBFs) [Ames2019CBFtheoryandapplications] is one of the promising approaches, ensuring forward invariance of a given set. CBFs can be easily implemented with quadratic programming and have been widely used in robotics applications [dawson2022barrier][Mitsioni2020][ykgpracbf]. Although CBFs accomplish the forward invariance of a given static constraint set, they are not applicable to applications where a constraint set is changed over time: a time-varying set. In this vein, using Time-Varying Control Barrier Functions (TVCBFs) is one of the potential ways to enforce time-varying constraints and successfully applied to human-assist wheelchair control [TvCBF2019] and dynamic obstacle avoidance [Huang2023].
However, since CBF-based controllers are highly dependent on the model accuracy, the forward invariance of a given set might not be satisfied if the model uncertainties such as parametric uncertainty and input disturbances are not taken into account. To resolve this issue, adaptive and robust control schemes have been recently proposed. Robust adaptive CBFs (RaCBFs) [lopez2020robust] introduce the tightened safe set by imposing robustness against the uncertainty and can be integrated with Set Membership IDentification (SMID) [SMID1992] to estimate the maximum bound of the uncertainty, so that less conservatism can be achieved. Another uncertainty, input disturbances are frequent in real-world implementations; these deteriorate the performance of CBFs-based controllers. The notion of Input-to-State Safety (ISSf) is defined to treat input disturbances, making a given safe set inflated based on the magnitude of the disturbances [ShishirISSfCBFs2019]; thus they enable minimizing safety violations. Notably, in [TEZUKA202244], TVCBFs are combined with ISSf conditions to impose robustness against the bounded input disturbances. However, they do not consider inherent model uncertainty in the system model like RaCBFs, so the violation of constraints could occur.
Figure 1: Block diagram of the control system including a robust adaptive time-varying control barrier function with a given time-varying safe set, $\mathcal{S}(t)$. Given an uncertain system with parametric uncertainty and additive disturbances, the safety filter ensures that the system states stay in a robust subset, $\mathcal{S}_{\bm{\theta}_{\delta}}^{r}\subseteq\mathcal{S}(t)$. Data-driven update method, Set Membership IDentification (SMID) is used to reduce the maximum possible error $|\bm{\tilde{\vartheta}}$$|$.
@@
We demonstrate the proposed method on robotic surface treatment application in this paper. Generally, the robotic surface treatment can be accomplished by maintaining a constant contact force between the polishing tool and the surface of a given workpiece [Li]. Many force control schemes including passive [Wang][Han] and active compliance controllers [Buckmaster] have been used in robotic surface treatment.
Even if the current robotic systems for surface treatment have focused on controlling constant contact forces to achieve an acceptable quality, they often neglect contact pressures over the surface, which leads to failure to keep uniform quality. Moreover, formal quality guarantees have not been fully investigated yet, which is a main challenge in robotic surface treatment applications [YkCASE2022][XIAO2020188]. To remedy this challenge, the proposed method is suitable for the robotic surface treatment application to accomplish quality guarantees based on a set invariance scheme. This is achieved by the fact that the proposed method regulates the time-varying reference contact force trajectory, which in turn governs the contact pressure, and then we ultimately regulate the quality of robotic surface treatment.
@@ II Motivating Example
In this section, we describe an application example that motivates this work. The proposed method is inspired by a robotic surface treatment task where the quality can be determined based on the material removal rate (MRR) [Li]. The MRR is modeled using Preston’s equation [Preston]:
@@ III-A Time-Varying Control Barrier Functions
where $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state, $f,g$ are locally Lipschitz continuous functions, $\bm{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ represents a control input. Consider a time-varying safety set $\mathcal{S}$ defined as $\mathcal{S}=\{\bm{x}\in\mathcal{X}\text{ }|\text{ }h(\bm{x},t)\geq 0\}$. A continuously differentiable function $h:\mathcal{X}\times\mathbb{R}\rightarrow\mathbb{R}$ is a Time-Varying Control Barrier Function (TVCBF) if there exist constants $K>0$ and $C>0$ satisfying the following for all $t\geq 0$ [TvCBF2019]:
@@ III-B Input-to-State Safety Control Barrier Functions
[…]
@@
where $\mathcal{S}\subseteq\mathcal{S}_{\delta}$ and $h_{\delta}(\bm{x})=h(\bm{x})+\gamma(\delta)$. We say that the system (5) is Input-to-State Safe (ISSf) if there exists $\gamma$ and $\delta$ such that for all $\bm{d}(t)$, $\mathcal{S}_{\delta}$ is forward invariant [ShishirISSfCBFs2019]. An ISSf control barrier function is defined as follows:
Theorem 1 ([ShishirISSfCBFs2019])
@@ III-C Robust adaptive Control Barrier Functions (RaCBFs)
In the presence of parametric uncertainty in the system, we consider the following control affine system and adaptive parameter estimator dynamics simultaneously [Taylor2020]:
[…]
@@
[…]
@@ III-D Set Membership IDentification (SMID)
Set Membership IDentification (SMID) is leveraged to compute a bound on the parameter estimation error based on input-output data [SMID1992]. We consider a linear regression equation
@@
where $d_{j}$ is a disturbance, $\bm{\theta}$ is a vector of unknown constant parameters, and $\mathcal{Y}_{j}$, $\mathcal{D}_{j}$ are measured quantities. Let $\mathcal{H}=\{\mathcal{Y}_{j},\mathcal{D}_{j}\}^{M}_{j=1}$ be a data history with size $M\in\mathbb{N}$. To bound the value of $\bm{\theta}^{*}$ in (8), we define $\mathcal{Y}_{j}=\dot{\bm{x}}(t_{j})-f(\bm{x}(t_{j}))-g(\bm{x}(t_{j}))\bm{u}(t_{j})$ and $\mathcal{D}_{j}=F(x(t_{j}))$, where $t_{j}$ is the time at which the data is recorded. Consider the following sequence of bounded parameter sets with $\Xi_{0}=[\underline{\bm{\theta}}_{0},\overline{\bm{\theta}}_{0}]$ and $\forall j\in M$:
@@
where $\underline{\bm{\theta}}_{0},\overline{\bm{\theta}}_{0}]$ is the initial lower and upper bound of the unknown parameter, and $\bm{1}_{n}$ is an $n$-dimensional vector of ones, and $\varepsilon\in\mathbb{R}$ is the precision variable that must be chosen such that $d_{j}||_{\infty}\leq\epsilon$ for all $j$. The sequence of sets, $\Xi_{t_{j}}$ can be calculated based on Linear Programming (LP) such that $\Xi_{t_{j}}\subseteq\Xi_{t_{j-1}}\subseteq\cdots\subseteq\Xi_{0}$ and $\bm{\theta}\in\Xi_{t_{j}}$ [smidbook].
@@ IV-A Parametric Adaptive System Model with Uncertainties
where $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state, $f,g$ are locally Lipschitz continuous functions, $\bm{u}\in\mathcal{U}\subset\mathbb{R}^{m}$ represents a control input, and $\bm{d}(t)\in\mathbb{R}^{m}$ is a bounded input disturbance, $\bm{d}(t)||_{\infty}\leq\delta\in\mathbb{R}_{+}$, and $F:\mathcal{X}\rightarrow\mathbb{R}^{n\times k}$ is smooth on $\mathcal{X}$ and models parametric uncertainty for $f$ such that $F(\bm{0})=\bm{0}$, and $\bm{\theta}^{*}\in\Theta\subset\mathbb{R}^{k}$ is the vector of unknown parameters. In the following, we include the time derivative term of $h_{r}$ to deal with time-varying constraints.
@@ IV-B Robust Adaptive Time-Varying Control Barrier Function
[…]
@@
A continuously differentiable function $h$ is a Robust adaptive Time-Varying Control Barrier Function (RaTVCBF) for the system (15) with bounded disturbance, $\bm{d}(t)||_{\infty}\leq\delta$ if there exist extended class $\mathcal{K}_{\infty}$ functions $\alpha$ and $\gamma$ such that for any $\bm{\theta}\in\Theta$ and for all $\bm{x}\in\mathcal{S}^{r}_{\bm{\theta}_{\delta}}$:
@@
[…]
@@
[…]
@@
The proof follows [lopez2020robust]. Let a candidate control barrier function $h$ be:
@@
$\dot{h}(\bm{x},\hat{\bm{\theta}},t)=\dot{h}_{r}(\bm{x},\hat{\bm{\theta}},t)+\tilde{\bm{\theta}}^{\top}\Gamma^{-1}\dot{\hat{\bm{\theta}}}\\$ (21)
@@
$).\frac{\partial h_{r}}{\partial\bm{x}}F(\bm{x})\Bigg(\hat{\bm{\theta}}-\Gamma\bigg(\frac{\partial h_{r}}{\partial\hat{\bm{\theta}}}\bigg$$^{\top}\Bigg)$. (22)
Then, we obtain the following equation and inequalities with $\lambda(\bm{x},\hat{\bm{\theta}},t)\triangleq{\hat{\bm{\theta}}}-\Gamma\big(\frac{\partial h_{r}}{\partial\hat{\bm{\theta}}}(\bm{x},\hat{\bm{\theta}},t)\big)^{\top}$ from [Taylor2020]:
@@
[…]
In the last inequality of (23), we choose $\tau(\bm{x},\hat{\bm{\theta}},t)=-\Big(\frac{\partial h_{r}}{\partial\bm{x}}F(\bm{x})\Big)^{\top}$ similar to (9) provided by [Taylor2020], and then the second term will be zero. Afterwards, we introduce the maximum possible error, $\tilde{\bm{\vartheta}}$, and if the following inequality is satisfied:
@@
[…]
@@
$\bm{u}_{\mathsf{safe}}=\quad\operatorname*{arg\,min}_{\bm{u}\in\mathcal{U}}\frac{1}{2}{||\bm{u}-\bm{u}_{\mathsf{nominal}}||}^{2}\quad\textbf{(RaTVCBF-QP)}$
$\text{s.t. }L_{f}h_{r}+L_{F}h_{r}\lambda(\bm{x},\hat{\bm{\theta}},t)+L_{g}h_{r}\bm{u}\geq$
@@ V Application to Robotic Surface Treatment
In this section, it is shown that the proposed method ensures that MRR (1) remains within an acceptable range across the entire surface. According to [Li], controlling MRR within the desired range ultimately achieves the desired quality guarantees. We verify the proposed method with numerical simulation on MATLAB and conduct experimental validation in a real robotic setup.
@@ V-A Robotic Surface Treatment
To ensure quality guarantees of the surface treatment via MRR as we mentioned in Section II, we enforce contact force constraints in (25). First, we define the parametric contact force, $\bm{f}_{c}\in\mathbb{R}^{n}$ from the Kelvin-Voigt model [kelvinModel] as:
@@
Figure 3: Pipeline of the robotic surface treatment. First, a user defines the desired MRR. Subsequently, in the force reference generator, we estimate contact area between a tool and a given surface and then calculate the desired contact pressure based on the desired MRR. Afterwards, the desired force is obtained from the desired contact pressure by using (1). Finally, we can determine force bounds (e.g. $\pm 15\%$ of the desired force) that are used as constraints to ensure the desired MRR in the dotted box.